1.     Find a point-normal form of the equation of the plane passing through P = (1, 2,-3) and having n =< 2,-1, 2 > as a normal.

2.     Determine in each case whether the given planes are parallel or perpendicular:

a)   x + y + 3z + 10 = 0 and x + 2y - z = 1

b)   3x - 2y + z - 6 = 0 and 4x + 2y - 4z = 0

a)     3x + y + z - 1 = 0 and - x + 2y + z + 3 = 0

b)     x - 3y + z + 1 = 0 and 3x - 4y + z - 1 = 0


1
Expert's answer
2021-06-28T16:36:34-0400

1. A point-normal form of the equation of the plane is


"2(x-1)-(y-2)+2(z+3)=0"

Or

"2x-y+2z+6=0"

2.

a) x+y+3z+10=0 and x+2y-z=1


"\vec n_1=\langle1,1,3\rangle, \vec n_2=\langle1,2,-1\rangle""\dfrac{1}{1}\not=\dfrac{2}{1}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.



"\vec n_1\cdot\vec n _2=1(1)+1(2)+3(-1)=0"

The vectors "\vec n_1" and "\vec n_2" are orthogonal.

Hence, the given planes are perpendicular.


b) 3x-2y+z-6=0 and 4x+2y-4z=0


"\vec n_1=\langle3,-2,1\rangle, \vec n_2=\langle4,2,-4\rangle""\dfrac{4}{3}\not=\dfrac{2}{-2}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.



"\vec n_1\cdot\vec n _2=3(4)+(-2)(2)+1(-4)=4\not=0"

The vectors "\vec n_1" and "\vec n_2" are not orthogonal.

Hence, the given planes are neither parallel nor perpendicular.


c) 3x+y+z-1=0 and -x+2y+z+3=0


"\vec n_1=\langle3,1,1\rangle, \vec n_2=\langle-1,2,1\rangle""\dfrac{-1}{3}\not=\dfrac{2}{1}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.



"\vec n_1\cdot\vec n _2=3(-1)+1(2)+1(1)=0"

The vectors "\vec n_1" and "\vec n_2" are orthogonal.

Hence, the given planes are perpendicular.


d) x-3y+z+1=0 and 3x-4y+z-1=0


"\vec n_1=\langle1,-3,1\rangle, \vec n_2=\langle3,-4,1\rangle""\dfrac{3}{1}\not=\dfrac{-4}{-3}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.



"\vec n_1\cdot\vec n _2=1(3)+(-3)(-4)+1(1)=16\not=0"

The vectors "\vec n_1" and "\vec n_2" are not orthogonal.

Hence, the given planes are neither parallel nor perpendicular.